OpenStudy (anonymous):

Can some one please help me on Simplifying the radical expression. Equation below in comments. :)

OpenStudy (anonymous):

\[\sqrt{14q} \times 2 \sqrt{4q}\]

OpenStudy (anonymous):

Jeez you guys never help me. This site is useless